Can someone help me with these 2.

All my answers are apparently wrong on number 11.

9A student stands at the edge of a cliff and throws a stone horizontally over the edge with a speed of v = 13.5 m/s. The cliff(d) Write the equations for the position of the stone with time, using the coordinates in the fiqure. (Use the following as n

11From the window of a building, a ball is tossed from a height y above the ground with an initial velocity of 8.10 m/s and ang
Answer #1

9.(a) co-ordinates of initial point are C = Om y = 36m origin Bottom of the ceild (6) Horizontal component (Vox), Vertical cobut Y= h-yl {as origin is at e bottom.} Y= h - 12gt2 y = 36.0-4.9+ using equation h=0+ $8+ Sto 2.71056] to 2.710 Sec (f) verte = tanpa so 0 = 63.091° 11. S-IMA 16 t: 4 sec. TITTT. (a) Initial co-ordinates of coin C: O y=% (b) Initial components of ve20= 7.786.t y= 2.232t+4xgxt? Height from which it was thrown. 8 = 2.232/4 + $*9.8x4²=165.728m y = h-y 1 y = 165.73 - 2.327-4

Sir feel free to ask if any doubt in comment section.

Please dont't forget to press like button if you liked this answer...Because it'll help me to improve my CF Score.

Thank you sir

Have a nice day.